# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 7. Chern forms of qas bundles
Let $(E,h)$ be a qas bundle, $h_{0}$ a smooth reference metric, and let $\hat{\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0})-I$. In this section, we will mostly drop $h$ and $h_{0}$ from the notation. In view of Theorem 6.5 (ii), $\hat{\bf s}(E):={\bf s}(E)-I={\bf s}_{1}(E)+{\bf s}_{2}(E)+\cdots$ is an endomorphism on ${\mathcal{QZ}}(X)$ of positive bidegree. Let $\hat{\bf s}(E)^{k}$ the composition of it $k$ times. We define the Chern endomorphism
@@
Just as for the Segre form, the Chern endomorphism yields a Chern form $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1\in{\mathcal{QZ}}(X)$. If $h$ is smooth, then $c(E,h,h_{0})$ equals the usual Chern form of $h$, cf. (2.4).
@@
(i) ${\bf c}(E)\wedge{\bf s}(E)={\bf s}(E)\wedge{\bf c}(E)=I$.
The statements in Proposition 6.4 and (ii) to (vi) of Theorem 6.5 hold if ${\bf s}(E)$ is replaced by ${\bf c}(E)$.
@@
Clearly ${\bf s}(E)$ and $\hat{\bf s}(E)$ commute, and hence the computation
${\bf s}(E)\wedge{\bf c}(E)=(I+\hat{\bf s}(E))\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E)^{k}=I,$
is legitimate, so that ${\bf s}(E)\wedge{\bf c}(E)={\bf c}(E)\wedge{\bf s}(E)=I$. The other equality in (i) follows in the same way.
The corresponding version of Proposition 6.4 with ${\bf s}(E)$ replaced by ${\bf c}(E)$ follows from Proposition 6.4 and Proposition 6.11.
It remains to prove the corresponding statements of Theorem 6.5 when ${\bf s}(E)$ is replaced by ${\bf c}(E)$. Part (ii) is obvious, and part (iii) for ${\bf c}(E)$ follows from (iii) for ${\bf s}(E)$.
To see (iv), notice that ${\bf s}(E,h,h_{0})-{\bf s}(E,h_{0})=\hat{\bf s}(E,h,h_{0})-\hat{\bf s}(E,h_{0})$ and that $\hat{\bf s}(E,h,h_{0})$ and $\hat{\bf s}(E,h_{0})$ commute. By Theorem 6.5 (iv), for each $k$, there are $w_{k}$ such that
@@
If $w=-w_{1}+w_{2}-\cdots$, then $dd^{c}w=(s(E,h,h_{0})-s(E,h_{0}))\wedge\mu$.
We get (v) for ${\bf c}$ from the statement for ${\bf s}$ by a formal computation. For instance, by (i) and (v) for ${\bf s}$, $I={\bf s}(E^{\prime}\oplus E)\wedge{\bf c}(E^{\prime}\oplus E)=s(E^{\prime})\wedge{\bf s}(E)\wedge{\bf c}(E^{\prime}\oplus E)$. Multiplying by ${\bf c}(E)$ and ${\bf c}(E^{\prime})$ from the left we get ${\bf c}(E^{\prime})\wedge{\bf c}(E)={\bf c}(E^{\prime}\oplus E)$. Finally, we get (vi) for ${\bf c}$ by repeated use of (vi) for ${\bf s}$. ∎
We do not know if in general, $c_{k}(E)=0$ for $k>{\rm rank\,}E$, as it is in the smooth case.
7.1. Chern forms for qas metrics with empty unbounded locus
Let us now assume that $h$ is a qas metric with empty unbounded locus $Z$. If $Y$ is as in the end of Example 6.10, then just as in the proof of Proposition 6.11, one has a well-defined operator
(7.2) ${\bf s}_{\ell_{k}}(E,h)\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h)1=\Pi_{*}\big{(}(dd^{c}\pi_{k}^{*}\psi_{k})^{\ell_{k}+r-1}\wedge\cdots\wedge\pi_{1}^{*}(dd^{c}\psi_{1})^{\ell_{1}+r-1}\big{)},$
which is commuting in the factors, since it is the direct image of a Bedford-Taylor product on ${\mathbb{P}}(E)\times_{X}\cdots\times_{X}{\mathbb{P}}(E)$. In particular, one may define products of currents like $s_{\ell_{k}}(E,h)\wedge\cdots\wedge s_{\ell_{1}}(E,h)$. Therefore, there is a well-defined current
$c(E,h)=\frac{1}{s(E,h)}=1+(s_{1}(E,h)^{2}-s_{2}(E,h))+\cdots$
just as is in the smooth case. If $h_{\epsilon}$ is the regularization from Proposition 6.4, then by Proposition 6.11, $c(E,h)=\lim c(E,h_{\epsilon})$. In particular,
@@ 7.1. Chern forms for qas metrics with empty unbounded locus
since it holds for $c_{k}(E,h_{\epsilon})$.
We do not know whether it holds for a general qas metric $h$, i.e., when $h$ is not assumed to be locally bounded, that $c_{k}(E,h,h_{0})=0$ for $k>{\rm rank\,}E$.
@@ 7.2. An alternative endomorphism
As will be discussed in the final section, there does not appear to exist a canonical choice of Segre or Chern forms associated to a qas metric, except for in low degrees. In this subsection, we study another choice of an extension of a Segre form as an endomorphism of ${\mathcal{QZ}}(X)$, which in turn leads to a different choice of Chern form. This Chern form is the one which appears in [1].
Assume that $E$ is equipped with a qas metric $h$, and a reference metric $h_{0}$. In this section, we will mostly drop $h,h_{0}$ from the notation of $s(E,h,h_{0})$, and we will let $E_{0}$ denote the bundle $E$ equipped with the metric $h_{0}$, and hence write $s(E_{0}):=s(E,h_{0})$.
Let us introduce the endomorphism
@@
If $Z=X$, then it is just ${\bf s}(E_{0})$ so it is of interest only when $Z$ has positive codimension, which we will assume from now on. Note that if we let $s(E,Z):={\bf s}(E,Z)1$, then
@@
The equation in (i) follows readily by definition of ${\bf s}(E,Z)$.
To prove (ii), note first that since ${\bf 1}_{Z}\mu$ is in ${\mathcal{QZ}}(X)$ and has support on $Z$, by Proposition 4.4, there is a current $\mu_{Z}\in{\mathcal{QZ}}(Z)$ such that ${\bf 1}_{Z}\mu=i_{*}\mu_{Z}$, where $i:Z\to X$ denotes the inclusion. By Theorem 6.5 (iv), there is $w^{\prime}\in{\mathcal{QZ}}(Z)$ such that
@@
To be precise, we have formulated Theorem 6.5 only for a complex manifold, but when want to apply it here on $Z$ which might be singular, i.e., a reduced analytic space. However, one may note either that the same result holds on reduced analytic spaces, with the same proof, or one might reduce to the case of manifolds by taking a resolution of singularities $\tau\colon Z^{\prime}\to Z$, finding the appropriate current on $Z^{\prime}$ and taking the push-forward to $Z$. Hence, (7.7) follows by (i) together with Theorem 6.5 (iii) by taking $w=i_{*}w^{\prime}$. ∎
@@
(7.10) ${\bf c}(E,Z)=\big{(}{\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E)\big{)}\wedge{\bf 1}_{X\setminus Z}\mu+{\bf c}(E_{0})\wedge{\bf 1}_{Z}\mu.$
By definition both $c(E)$ and $c(E,Z)$ are equal to the usual Chern form in $X\setminus Z$ if the metric on $E$ is smooth there. However, in general is $c(E,Z)\neq c(E)$ across $Z$, see Example 8.2.
@@
Part (i) follows in the same way as (i) in Proposition 7.1.
We may decompose
@@
By (7.7), for each $\ell$, there exists $w_{\ell}$, which is the push-forward of a current on $Z$, such that
@@
so we may take
$w=\sum s(E_{0})^{k-\ell-1}\wedge w_{\ell}$
to prove (ii).
For (iii), recall that $M^{E}\wedge\mu:={\bf 1}_{Z}{\bf s}(E)\wedge\mu$ and ${\bf s}^{\prime}(E)\wedge\mu={\bf 1}_{X\setminus Z}{\bf s}(E)\wedge\mu.$ We will prove by induction that
@@
when acting on ${\bf 1}_{X\setminus Z}\mu$. By definition it holds for $k=1$. Assume that it holds for $k$. When applying ${\bf s}(E)\wedge{\bf 1}_{X\setminus Z}+{\bf s}(E_{0})\wedge{\bf 1}_{Z}$ to the right hand side of (7.11) the first mapping only acts on the first term and gives $\hat{\bf s}(E)\wedge\hat{\bf s}^{\prime}(E)^{k}=\hat{\bf s}^{\prime}(E)^{k+1}+M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k}$. The second mapping neglects the first term and just gives another factor $s(E_{0})$ on the terms in the sum. Thus (7.11) holds. Thus
${\bf c}(E,Z)=I+\sum_{k=1}^{\infty}(-1)^{k}\Big{(}\hat{\bf s}^{\prime}(E)^{k}+\sum_{\ell=0}^{k-1}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}\Big{)}=\\ {\bf c}^{\prime}(E)+\sum_{\ell=0}^{\infty}\sum_{k=\ell+1}^{\infty}(-1)^{k}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}=\\ {\bf c}^{\prime}(E)-\sum_{\ell=0}^{\infty}(-1)^{\ell}{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf c}^{\prime}(E)={\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E).$
@@
The action on ${\bf 1}_{Z}\mu$ is just $\hat{\bf s}(E_{0})^{k}\wedge{\bf 1}_{Z}\mu$. Thus (iii) follows. ∎
Remark 7.5.
One can check directly that $c(E,Z)={\bf c}(E,Z)1$ is cohomologous with $c(E_{0})$. In fact, accepting that the operations are legitimate we have, if
$dd^{c}u=s(E)-s(E_{0})=M^{E}+s^{\prime}(E)-s(E_{0}),$
that
$-dd^{c}(c(E_{0})\wedge u\wedge c^{\prime}(E))=-c(E_{0})\wedge(s(E)-s(E_{0}))\wedge c^{\prime}(E)=c^{\prime}(E)-c(E_{0})\wedge M^{E}\wedge c^{\prime}(E)-c(E_{0}).$
8. Examples and remarks
In this final section, we discuss to which extent there is a canonical choice of $s(E,h)$ (and other characteristic forms, i.e., Chern forms $c(E,h)$) when $h$ is a qas metric. In case the unbounded locus $Z$ of $h$ is empty, there seems to be a canonical choice of Segre form $s(E,h)$, given, e.g., by Segre forms $s(E,h_{\epsilon})$ of appropriate decreasing regularizations $h_{\epsilon}$ of $h$ as ${\epsilon}\to 0$, cf. Proposition 6.4. When $Z\neq\emptyset$, the form $s(E,h,h_{0})$ will typically depend on $h_{0}$, see [17, Example 8.2]. In particular, it follows by Theorem 1.4 (i) that limits $s(E,h_{\epsilon})$ as ${\epsilon}\to 0$ might depend on the sequence $h_{\epsilon}$, even when $h_{\epsilon}$ is a sequence of Griffiths negative smooth metrics decreasing to $h$. Hence, there is in general no canonical choice of a Segre form $s(E,h)$ depending only on $h$.
To be a bit more precise, it follows from our results that in low enough degrees, there is a natural canonical choice of $s_{k}(E,h)$. Let $Z$ denote the set where $h$ is not locally bounded, and let $p={\rm codim\,}Z$. As long as $k\leq p$, $s_{k}(E,h,h_{0})$ is independent of $h_{0}$ by Remark 6.6 (i). On the other hand, [17, Example 8.2] mentioned above shows that $s_{k}(E,h,h_{0})$ might depend on $h_{0}$ for any $k>p$.
In case $k<p$, any definition of $s_{k}(E,h)$ as a closed current of order 00 is in fact uniquely determined by its restriction to $X\setminus Z$. Indeed, because the difference between two such choices is a normal $(k,k)$-current with support on an analytic subset of codimension $>k$, it must vanish, cf. e.g., [13, Corollary III.2.11]. In particular, $s_{k}(E,h)=s_{k}^{\prime}(E,h)$, which coincides with the currents considered by Xia (for more general classes of metrics), [21], based on using the non-pluripolar Monge-Ampère product.
In the case when $k=p$, one may by Theorem 1.4 and Proposition 6.4 find locally an appropriate decreasing sequence of smooth metrics $h_{\epsilon}$ such that $s_{p}(E,h_{\epsilon})\to s_{p}(E,h)$, independent of the choice of such a sequence. If the metric $h$ is furthermore Griffiths negative, then by [17], $s_{k}(E,h)$ may be obtained as $\lim_{{\epsilon}\to 0}s_{k}(E,h_{\epsilon})$ for any sequence of smooth Griffiths negative metrics $h_{\epsilon}$ decreasing to $h$ as ${\epsilon}\to 0$ for any $k\leq p$.
Similar statements as above will hold also for products

### v2

@@ 7. Chern forms of qas bundles
Let $(E,h)$ be a qas bundle, $h_{0}$ a smooth reference metric, and let $\hat{\bf s}(E,h,h_{0})={\bf s}(E,h,h_{0})-I$. In this section, we will sometimes drop $h$ and $h_{0}$ in the notation. In view of Theorem 6.7 (ii), $\hat{\bf s}(E):={\bf s}(E)-I={\bf s}_{1}(E)+{\bf s}_{2}(E)+\cdots$ is an endomorphism on ${\mathcal{QZ}}(X)$ of positive bidegree. Let $\hat{\bf s}(E)^{k}$ the composition of it $k$ times. We define the Chern endomorphism
@@
The statements in Theorem 6.5 and (ii)-(vi) of Theorem 6.7 hold if ${\bf s}(E,h,h_{0})$ is replaced by ${\bf c}(E,h,h_{0})$.
@@
Clearly ${\bf s}(E,h,h_{0})$ and $\hat{\bf s}(E,h,h_{0})$ commute, and hence the computation
${\bf s}(E,h,h_{0})\wedge{\bf c}(E,h,h_{0})=(I+\hat{\bf s}(E,h,h_{0}))\sum_{k=0}^{\infty}(-1)^{k}\hat{\bf s}(E,h,h_{0})^{k}=I,$
is legitimate. The other equality in (7.2) follows in the same way.
We now consider the analogue for $\bf c$} of Theorem 6.5, so assume that $h$ is qas and non-denegerate. Then clearly $c(E,h)$ is independent of $h_{0}$. If $h_{\epsilon}$ is a local regularization, then it follows from (6.17) that the analogue for ${\bf c}(E,h)$ of part (ii) holds. Part (iii) implies its analogue for $\bf c$}: Suppressing $h$ and $h^{\prime}$ in the notation, by (i) and (iii) for $\bf s$} we have $I={\bf s}(E^{\prime}\oplus E)\wedge{\bf c}(E^{\prime}\oplus E)={\bf s}(E)\wedge{\bf s}(E^{\prime})\wedge{\bf c}(E^{\prime}\oplus E)$. Multiplying by ${\bf c}(E^{\prime})$ from the right and ${\bf c}(E)$ from the left we get ${\bf c}(E)\wedge{\bf c}(E^{\prime})={\bf c}(E^{\prime}\oplus E)$.
From (2.4) we see that ${\bf c}_{k}(E,h)$ can be expressed as a sum of commuting compositions ${\bf s}_{\ell_{1}}(E,h)\cdots{\bf s}_{\ell_{m}}(E,h)$. That is,
${\bf c}_{1}(E,h)=-{\bf s}_{1}(E,h),\quad{\bf c}_{2}(E,h)={\bf s}_{2}(E,h)-{\bf s}_{1}(E,h)^{2},\quad etc.$
Now (iv) follows for $\bf c$}.
It remains to prove the analogous statements of Theorem 6.7 when ${\bf s}(E)$ is replaced by ${\bf c}(E)$. Part (ii) is obvious, and part (iii) for ${\bf c}(E)$ follows from (iii) for ${\bf s}(E)$. To see (iv), notice that ${\bf s}(E,h,h_{0})-{\bf s}(E,h_{0})=\hat{\bf s}(E,h,h_{0})-\hat{\bf s}(E,h_{0})$ and that $\hat{\bf s}(E,h,h_{0})$ and $\hat{\bf s}(E,h_{0})$ commute. By Theorem 6.7 (iv), for each $k$, there are $w_{k}$ such that
@@
If $w=-w_{1}+w_{2}$-$\cdots$, then $dd^{c}w=({\bf c}(E,h,h_{0})-{\bf c}(E,h_{0}))\wedge\mu$. We get (v) for $\bf c$} from the statement for $\bf s$} by the same formal computation as above in the non-degenerate case. Finally, we get (vi) for $\bf c$} by repeated use of (vi) for $\bf s$}. ∎
Proof of Theorem 1.5.
The theorem follows from Proposition 7.1 and the fact that the Chern form $c(E,h,h_{0}):={\bf c}(E,h,h_{0})1$ is in ${\mathcal{QZ}}(X)$, cf. Section 4 in the same way that Theorem 1.3 follows from Theorem 6.7 and the fact that the Segre form $s(E,h,h_{0})$ is in ${\mathcal{QZ}}(X)$. ∎
If $h$ is qas and non-degenerate, then by the analogue of Theorem 6.5 (ii),
$c(E,h)=\lim_{{\epsilon}\to 0}c(E,h_{\epsilon})$
where $h_{\epsilon}$ are smooth metrics. In particular,
@@
since this holds for $c_{k}(E,h_{\epsilon})$. We do not know whether $c_{k}(E,h,h_{0})=0$ for $k>{\rm rank\,}E$ if $h$ is a general qas metric.
@@ 7.1. An alternative endomorphism
Let $(E,h)$ be a qas bundle on $X$ and a let $h_{0}$ be a smooth reference metric. We shall now consider an alternative extension to an endomorphism on ${\mathcal{QZ}}(X)$ which leads to a different Chern form, that appeared in [1]. We will use the notation $s(E):=s(E,h,h_{0})$ and $s(E_{0})=s(E,h_{0})$. Let $Z$ be the degeneracy locus of $h$.
We introduce the endomorphism
@@
If $Z=X$, then it is just ${\bf s}(E_{0})$ so it is of interest only when $Z$ has positive codimension, which we will assume from now on. Note that if $s(E,Z):={\bf s}(E,Z)1$, then
@@
Equation (7.6) follows directly from the definition of ${\bf s}(E,Z)$. To prove (ii), first notice that since ${\bf 1}_{Z}\mu$ is in ${\mathcal{QZ}}(X)$ and has support on $Z$, by Proposition 4.4, there is a current $\mu_{Z}\in{\mathcal{QZ}}(Z)$ such that ${\bf 1}_{Z}\mu=i_{*}\mu_{Z}$, where $i\colon Z\to X$ denotes the inclusion. By Theorem 6.7 (iv) there is $w^{\prime}\in{\mathcal{QZ}}(Z)$ such that
@@
To be precise, we have formulated Theorem 6.7 only for a complex manifold, but the proof if parts (iii) and (iv) works just as well if $Z$ is any reduced analytic space. Hence, (7.7) follows by (i) together with Theorem 6.7 (iii) with $w=i_{*}w^{\prime}$. ∎
@@
(7.10) ${\bf c}(E,Z)\wedge\mu=\big({\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E)\big)\wedge{\bf 1}_{X\setminus Z}\mu+{\bf c}(E_{0})\wedge{\bf 1}_{Z}\mu.$
By (ii), $c(E,Z):={\bf c}(E,Z)\wedge 1$ is cohomologous with $c(E)$ and hence it defines the expected Bott-Chern class.
By definition both $c(E)$ and $c(E,Z)$ are equal to the usual Chern form in $X\setminus Z$ if the metric on $E$ is smooth there. However, in general $c(E,Z)\neq c(E)$ across $Z$, see Example 8.1.
@@
Part (i) follows in the same way as (i) in Proposition 7.1. To see (ii) we first write
@@
By (7.7), for each $\ell$ there exists $w_{\ell}$, which is the push-forward of a current on $Z$, such that
@@
so we get (ii) with
$w=\sum s(E_{0})^{k-\ell-1}\wedge w_{\ell}.$
For (iii), recall that $M^{E}\wedge\mu:={\bf 1}_{Z}{\bf s}(E)\wedge\mu$ and ${\bf s}^{\prime}(E)\wedge\mu={\bf 1}_{X\setminus Z}{\bf s}(E)\wedge\mu.$ One checks that
(7.11) $\hat{\bf s}(E,Z)=\hat{\bf s}(E)\wedge{\bf 1}_{X\setminus Z}+\hat{s}(E_{0})\wedge{\bf 1}_{Z}$
We will prove by induction that
@@
when acting on ${\bf 1}_{X\setminus Z}\mu$. By definition it holds for $k=1$. Assume that it holds for $k$. When applying (7.11) to the right hand side of (7.12) the first mapping only acts on the first term and gives $\hat{\bf s}(E)\wedge\hat{\bf s}^{\prime}(E)^{k}=\hat{\bf s}^{\prime}(E)^{k+1}+M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k}$. The second mapping neglects the first term and just gives another factor $\hat{\bf s}(E_{0})$ on the terms in the sum. Thus (7.12) holds. Now
${\bf c}(E,Z)=I+\sum_{k=1}^{\infty}(-1)^{k}\Big(\hat{\bf s}^{\prime}(E)^{k}+\sum_{\ell=0}^{k-1}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}\Big)=\\ {\bf c}^{\prime}(E)+\sum_{\ell=0}^{\infty}\sum_{k=\ell+1}^{\infty}(-1)^{k}\hat{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf s}^{\prime}(E)^{k-1-\ell}=\\ {\bf c}^{\prime}(E)-\sum_{\ell=0}^{\infty}(-1)^{\ell}{\bf s}(E_{0})^{\ell}\wedge M^{E}\wedge\hat{\bf c}^{\prime}(E)={\bf c}^{\prime}(E)-{\bf c}(E_{0})\wedge M^{E}\wedge{\bf c}^{\prime}(E)$
@@
when acting on ${\bf 1}_{X\setminus Z}\mu$. The action on ${\bf 1}_{Z}\mu$ is just $\hat{\bf s}(E_{0})^{k}\wedge{\bf 1}_{Z}\mu$. Thus (iii) follows. ∎
8. Remarks and examples
If $(E,h)$ is qas and its degeneracy locus $Z$ is empty, then, in view of (6.18) and (6.19), the Segre form $s(E,h)$ defined here seems to be the only reasonable one and thus canonical.
If $Z\neq\emptyset$, then any closed current of order 0 that coincides with $s_{k}(E,h)$ on $X\setminus Z$ is unique if $k<{\kappa}:={\rm codim\,}Z$. This follows from the dimension principle for normal currents, see, e.g., [13, Corollary III.2.11]. In particular, $s_{k}(E,h)=s_{k}^{\prime}(E,h)$, which coincides with the currents considered by Xia (for more general classes of metrics), [21], based on the non-pluripolar Monge-Ampère product. Since the class of $s_{k}(E,h,h_{0})$ in ${\mathcal{QB}}(X)$ is independent of $h_{0}$ it follows from Proposition 4.15 that $s_{\kappa}(E,h,h_{0})$ is independent of $h_{0}$. If the metric $h$ in addition is Griffiths negative, then by [17], $s_{k}(E,h)$ may be obtained as $\lim_{{\epsilon}\to 0}s_{k}(E,h_{\epsilon})$ for any sequence of smooth Griffiths negative metrics $h_{\epsilon}$ decreasing to $h$ as ${\epsilon}\to 0$ for any $k\leq{\kappa}$. On the other hand, [17, Example 8.2] shows that $s_{k}(E,h,h_{0})$ might depend on $h_{0}$ for any $k>{\kappa}$. It follows from Theorem 1.3 (i) that the limits $s(E,h_{\epsilon})$ as ${\epsilon}\to 0$ might depend on the sequence $h_{\epsilon}$, even when $h_{\epsilon}$ is a sequence of Griffiths negative smooth metrics decreasing to $h$. Therefore it does not seem to be any ’canonical’ choice of Segre form $s_{k}(E,h)$ for $k>{\kappa}$.
Similar statements as above hold also for products